# Archilles’ Heel in Semi-open LLMs: Hiding Bottom against Recovery Attacks

### v1

@@ A.1 Model Overview
The recovered model $f(\mathbf{X};\bm{\theta})$ is structured as a sequence of $L$ transformer layers,
@@
where $Q_{i}\in\mathbb{R}^{d\times d_{Q}}$, $K_{i}\in\mathbb{R}^{d\times d_{Q}}$ represent projection parameter matrices. Here, the $\alpha L$-th layer is the recovered layer and the others are the public layers. For simplicity, we use the function $\hat{\varphi}_{\alpha L}$ to denote mapping of the recovered layer, i.e., $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$.
@@ A.2 Bounds on Different Orthogonal Components
The equation equation 6c is due to $\bm{u}_{l,k}$ are the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. The inequality equation 6e is because when $\bm{v}=\frac{\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}}{\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k}){u}_{l,k}\right\|_{2}}$, we have the maximum value.
@@
Let $\sigma_{i1},\sigma_{i2},\ldots,\sigma_{in}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Applying the Perron–Frobenius theorem for Markov matrices Lemmens & Nussbaum (2012), we deduce that for the matrix $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$, there exists only one eigenvalue equal to 1, while all other eigenvalues in absolute value are strictly less than 1. Without loss of generality, we assume $\sigma_{i1}=1$, implying $\left|\sigma_{ij}\right|<1$ for $j\neq 1$. Recalling the definition of $\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)$ and considering the linear operation, we can rewrite it as follows:
@@
where equation 11a is induced by substituting the equation equation 10 into $\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$, equation 11b is due to $\mathbb{I}_{n}^{\top}\bm{u_{j}}=0$ for all $j\neq 1$, equation 11c follows the fact that $\sigma_{i1}=1$
@@ A.3 Proof of Theorem 1
We first prove the following result. For simplicity of notations, we use $f(\mathbf{X})\left[p\right]$ to denote the $p$-th ($1\leq p\leq d$) column of the the recovered model $f(\mathbf{X})$, where the parameters in the $\alpha L$-th layer is replaced with the matrices $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$. We use the function $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$ to denote the mapping of the $(\alpha L)$-th layer. Then we are going to show that there exists $\alpha^{\star}=\log_{2}{\frac{2}{1+\beta_{D}}}$ and $10<\beta_{D}<$ 1 makes the following equations hold.
@@
(2) For any $\alpha>\alpha^{\star}$, there exists a sequence of matrix $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any recovered matrix $K_{\alpha L}$ and $Q_{\alpha L}$, we have $\left\|K_{i}\right\|_{2}\leq D,\left\|Q_{i}\right\|_{2}\leq D$, we have,
@@
Based on Lemma equation 1, we obtain that
@@
Based on Lemma equation 2, we know that
@@
We firstly prove the equation equation 12. When
@@
where the inequality equation 17c is based on the inequality equation 14 and equation 15. The inequality equation 17d is based on Lemma equation 3. Therefore, if $\alpha<\log_{2}{\frac{2}{1+\beta_{D}}}$ and $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|\neq 0$, then we have $\lim_{L\to\infty}\left(\frac{1+\beta_{D}}{2^{1-\alpha}}\right)^{L}=0$. Now we can consider when $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0.$ In fact, it is easy to show that this can only happens when $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ belong to certain sets making $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0,$ which corresponds to zero measure set $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ depending on the input $\mathbf{X}$. Since the input space is countable, therefore, the union $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$ are also zero-measure sets.
To prove equation equation 13, let $K^{\star}$, $Q^{\star}$ with $\left\|K^{\star}\right\|_{2}\leq D,\left\|Q^{\star}\right\|_{2}\leq D$ satisfy the following condition,
@@
Let $\bm{v}^{\star}$ be the solver of the above optimization problem equation 18 and consider the $K_{l}=K^{\star}$, $Q_{l}=Q^{\star}$ and $\mathbf{X}^{\star}=[\bm{v}^{\star},\bm{v}^{\star},\cdots,\bm{v}^{\star}]$. Clearly, $\bm{v}^{\star}\perp\mathbb{I}_{n}.$ Assume there exists $\bm{u}:\left\|\bm{u}^{\star}\right\|_{2}=1$ satisfying $\bm{u}^{\star}\perp\mathbb{I}_{n}$, $\bm{u}^{\star}\perp\bm{v}^{\star}$, therefore we can rewrite $f(\mathbf{X}^{\star})\left[p\right]$ as follows,
@@
For any $1\leq l\leq L$, based on Lemma equation 1, we know that
@@
where equation equation 23b is based on equation 19, equation equation 23e is based on equation 22 and equation 15. When $\alpha>\log_{2}{\frac{2}{1+\beta_{D}}}$, we have $\lim_{L\to\infty}\left(\frac{2^{1-\alpha}}{1+\beta_{D}}\right)^{L}=0$. Thus we have $\lim_{L\to\infty}\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star}\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=\sqrt{2}$. This indicates that the $p$-th column of the output matrix $f(\mathbf{X}^{\star})$ is not parallel to $\mathbf{I}_{n}$ for any $p$. This further indicates that the output matrix does not have the identical vector in each row. ∎
@@ A.4 Technical Lemma
To demonstrate equation 24, we equivalently show
@@
Clearly, equation 28 holds true for any $x\in(0,1)$. ∎
@@ Appendix B Experiment Details
The foundation models we use in our experiments are selected from open-source repositories, and Table 6 shows the basic information of the models and their sources. Specifically, we employ Llama2-70B-chat [^2] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf, Llama2-7B-chat [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, and Mistral-7B-v0.1444https://huggingface.co/mistralai/Mistral-7B-v0.1. For smaller models, we select Phi-2555https://huggingface.co/microsoft/phi-2 and Phi-1.5666https://huggingface.co/microsoft/phi-1_5. We also consider OPT model [^7] https://huggingface.co/facebook/opt-350m, which has only 350 million parameters and 24 decoder layers.
@@
B.2 Recovery Attacks.
Attack implementation details. In performing FT-all and FT-closed model recovery attacks, we adhere to the training hyper-parameters outlined in the Llama2 report (Touvron et al., 2023), employing the AdamW optimizer with a cosine learning rate scheduler. The initial learning rate is set to $2\times 10^{-5}$, with a weight decay of 0.1, a batch size of 128, and bfloat16 precision for input sequences of 512 tokens. The LLaMA2-70B model is trained for 3 epochs with a random seed of 42, while other models are trained for 5 epochs across three seeds: 42, 1234, and 20. Despite limiting training to 3 epochs for the 70B model, the training loss stabilized effectively. Our implementation builds upon the llama-recipes repository provided by META.
For SEM attacks, distinct configurations were employed for SCARA and SAP-DP. In the case of SCARA, embeddings from the closed-source components were collected and paired with the input data to train a substitute model. In contrast, for SAP-DP, embeddings from the sixth decoder layer and the model’s final logits were utilized to construct the training dataset. In accordance with (Tamber et al., 2024), we applied a learning rate of 1.5e-4, a weight decay of 0.01, and a linear learning rate scheduler with 500 warmup steps. Both training and validation batch sizes were set to 32, with MSE as the loss function. SCARA was trained for 30 epochs due to its smaller model size, whereas SAP-DP was trained for 5 epochs.
All recovery experiments were conducted on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, utilizing PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
Base 51k Recovery Dataset. We ensure dataset coverage and reliability by using a 1:1 ratio of the MMLU auxiliary training set  [^8] https://github.com/hendrycks/test and Alpaca dataset  [^9] https://github.com/tatsu-lab/stanford_alpaca/blob/main/alpaca_data.json, extracting 25.5k samples from each. From the MMLU auxiliary training data (Hendrycks et al., 2021), we sample 50%, and from Alpaca (Taori et al., 2023), we use a step size of 2 to enhance diversity. The datasets are then formatted for model training, applying Alpaca and MMLU prompts from Table 7.
@@
Extra Recovery Datasets. To enhance dataset diversity, the 100K, 200K, 300K, and 500K datasets integrate additional specialized sources. As detailed in Table 8, these sources include Baize (Xu et al., 2023) (158K English multi-turn conversations via ChatGPT’s self-chat), MathInstruct (Yue et al., 2023) (260K curated math instruction instances focusing on hybrid reasoning), and OpenOrca (Mukherjee et al., 2023) (augmented FLAN collection with 1M GPT-4 completions and 3.2M GPT-3.5 completions). These enrichments are intended to support complex computational and theoretical tasks, offering broader topic coverage.
Table 8: Composition of variously sized datasets
Raw Data Set 51k 100k 200k 300k 500k
Alpaca 25.5 50 40 50 50
MMLU auxiliary training set 25.5 50 40 100 100
Baize-MedQuAD 0 0 40 50 50
Baize-Quora 0 0 40 50 50
Baize-Stackoverflow 0 0 40 50 50
MathInstruct 0 0 4 6 20
OpenOrca 0 0 0 0 180
Validation Datasets. Table 9 outlines the composition of the validation datasets. For Validation Dataset 1, we extracted 50% from each of the 57 MMLU validation sub-datasets, totaling 1.5K instances, paired with Alpaca data selected using a step size of 751. This dataset is used with the 51K and 100K training sets. For larger training sets (200K, 300K, and 500K), Validation Dataset 2 was created by adding 400 instances from three Baize subsets, expanding the validation set to 4.0K.
@@
Baize-MedQuAD 0 850
Baize-Quora 0 850
Baize-Stackoverflow 0 850
Total Length 1516 4066
@@
In this section, we provide further details on the baselines used in our comparisons: SAP-DP and fully-closed. These schemes represent different strategies, each with distinct trade-offs in terms of customizability and resilience against model recovery attacks.
SAP. The Split-and-Privatize (SAP) framework (Shen et al., 2023) offers an approach to balance between protecting model privacy and data privacy while maintaining competitive performance. Specifically, the SAP framework keeps the bottom six encoder layers open, allowing user access and fine-tuning while closing the deeper layers on the vendor.
@@
Fully-closed. Following (Eiras et al., 2024), we use the fully-closed approach as a baseline. This assumes the adversary has no access to internal model parameters, treating the model as a black-box, where only output data can be collected. We slightly broaden this setup by assuming the adversary knows the model’s architecture but no other details. Thus, recovering the fully-closed model involves using the collected data to retrain a model with the same architecture to restore its general functionality.
B.4 Implementation Details of SCARA.
Evaluation Datasets. We created a 1.5K Evaluation Set to assess model resilience under various closed-sourcing strategies. This set includes 50% of entries from each of the 57 MMLU validation sub-datasets (Hendrycks et al., 2021), distinct from Validation Set outlined in Table 9. Additionally, we selected an equal number of Alpaca dataset (Taori et al., 2023), using a step size of 751, ensuring no overlap with the Validation Set.
Hyper-parameter Sensitivity. As shown in Figure 9, we evaluate SCARA’s sensitivity to tolerance magnitude $\varepsilon$, adjusting it from 0.05 to 1 in 0.05 increments while calculating the $\Delta\text{ARR}$ for six recovered models. The results indicate that SCARA is minimally sensitive to changes in $\varepsilon$, with $\Delta\text{ARR}$ values stabilizing as $\varepsilon$ increases. This stability arises from the need for a smaller closed-sourced layer at higher $\varepsilon$, allowing the condition $R(I)\leq(1+\varepsilon)R([L])$ to be met with fewer layers. Additionally, the increase in $\Delta\text{ARR}$ is smaller for larger models, suggesting that privatizing more parameters beyond a certain point offers diminishing returns in resilience.

### v2

@@ A.1 Model Overview
The distilled model $f(\mathbf{X};\bm{\theta})$ is structured as a sequence of $L$ transformer layers,
@@
where $Q_{i}\in\mathbb{R}^{d\times d_{Q}}$, $K_{i}\in\mathbb{R}^{d\times d_{Q}}$ represent projection parameter matrices. Here, the $\alpha L$-th layer is the distilled layer and the others are the public layers. For simplicity, we use the function $\hat{\varphi}_{\alpha L}$ to denote mapping of the distilled layer, i.e., $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$.
@@ A.2 Bounds on Different Orthogonal Components
The equation (5c) is due to $\bm{u}_{l,k}$ are the eigenvectors of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{l})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. The inequality (5e) is because when $\bm{v}=\frac{\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k})\bm{u}_{l,k}}{\left\|\sum_{k=2}^{n}\alpha_{pk}(1+\sigma_{l,k}){u}_{l,k}\right\|_{2}}$, we have the maximum value.
@@
Let $\sigma_{i1},\sigma_{i2},\ldots,\sigma_{in}$ denote the eigenvalues of $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{i}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$. Applying the Perron–Frobenius theorem for Markov matrices (Lemmens & Nussbaum, 2012), we deduce that for the matrix $\operatorname{softmax}\left(\frac{\mathbf{X}Q_{l}(\mathbf{X}K_{i})^{\top}}{\sqrt{d_{Q}}\|\mathbf{X}\|^{2}}\right)$, there exists only one eigenvalue equal to 1, while all other eigenvalues in absolute value are strictly less than 1. Without loss of generality, we assume $\sigma_{i1}=1$, implying $\left|\sigma_{ij}\right|<1$ for $j\neq 1$. Recalling the definition of $\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)$ and considering the linear operation, we can rewrite it as follows:
@@
where (10a) is induced by substituting the equation (9) into $\left|\mathbb{I}_{n}^{\top}\varphi_{i}\left(\mathbf{X};K_{i},Q_{i}\right)[p]\right|$, (10b) is due to $\mathbb{I}_{n}^{\top}\bm{u_{j}}=0$ for all $j\neq 1$, (10c) follows the fact that $\sigma_{i1}=1$
@@ A.3 Proof of Theorem 1
We first prove the following result. For simplicity of notations, we use $f(\mathbf{X})\left[p\right]$ to denote the $p$-th ($1\leq p\leq d$) column of the the distilled model $f(\mathbf{X})$, where the parameters in the $\alpha L$-th layer is replaced with the matrices $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$. We use the function $\hat{\varphi}_{\alpha L}(\mathbf{X})=\varphi_{\alpha L}(\mathbf{X};\hat{K}_{\alpha L},\hat{Q}_{\alpha L})$ to denote the mapping of the $(\alpha L)$-th layer. Then we are going to show that there exists $\alpha^{\star}=\log_{2}{\frac{2}{1+\beta_{D}}}$ and $10<\beta_{D}<$ 1 makes the following equations hold.
@@
(2) For any $\alpha>\alpha^{\star}$, there exists a sequence of matrix $\{K_{i},Q_{i}\}_{i\geq 1}$ such that for any distilled matrix $K_{\alpha L}$ and $Q_{\alpha L}$, we have $\left\|K_{i}\right\|_{2}\leq D,\left\|Q_{i}\right\|_{2}\leq D$, we have,
@@
Based on Lemma (1), we obtain that
@@
Based on Lemma (2), we know that
@@
We firstly prove the equation (11). When
@@
where the inequality (16c) is based on the inequality (13) and (14). The inequality (16d) is based on Lemma (3). Therefore, if $\alpha<\log_{2}{\frac{2}{1+\beta_{D}}}$ and $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|\neq 0$, then we have $\lim_{L\to\infty}\left(\frac{1+\beta_{D}}{2^{1-\alpha}}\right)^{L}=0$. Now we can consider when $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0.$ In fact, it is easy to show that this can only happens when $\hat{K}_{\alpha L}$ and $\hat{Q}_{\alpha L}$ belong to certain sets making $\left|\mathbb{I}_{n}^{\top}f(\mathbf{X})[p]\right|=0,$ which corresponds to zero measure set $\mathcal{K}(\mathbf{X})$ and $\mathcal{Q}(\mathbf{X})$ depending on the input $\mathbf{X}$. Since the input space is countable, therefore, the union $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{K}(\mathbf{X})$ and $\cup_{\mathbf{X}\in\mathcal{X}}\mathcal{Q}(\mathbf{X})$ are also zero-measure sets.
To prove equation (12), let $K^{\star}$, $Q^{\star}$ with $\left\|K^{\star}\right\|_{2}\leq D,\left\|Q^{\star}\right\|_{2}\leq D$ satisfy the following condition,
@@
Let $\bm{v}^{\star}$ be the solver of the above optimization problem (17) and consider the $K_{l}=K^{\star}$, $Q_{l}=Q^{\star}$ and $\mathbf{X}^{\star}=[\bm{v}^{\star},\bm{v}^{\star},\cdots,\bm{v}^{\star}]$. Clearly, $\bm{v}^{\star}\perp\mathbb{I}_{n}.$ Assume there exists $\bm{u}:\left\|\bm{u}^{\star}\right\|_{2}=1$ satisfying $\bm{u}^{\star}\perp\mathbb{I}_{n}$, $\bm{u}^{\star}\perp\bm{v}^{\star}$, therefore we can rewrite $f(\mathbf{X}^{\star})\left[p\right]$ as follows,
@@
For any $1\leq l\leq L$, based on Lemma (1), we know that
@@
where equation (22b) is based on (18), equation (22e) is based on (21) and (14). When $\alpha>\log_{2}{\frac{2}{1+\beta_{D}}}$, we have $\lim_{L\to\infty}\left(\frac{2^{1-\alpha}}{1+\beta_{D}}\right)^{L}=0$. Thus we have $\lim_{L\to\infty}\left\|\frac{f(\mathbf{X}^{\star})\left[p\right]}{\left\|f(\mathbf{X}^{\star}\left[p\right]\right\|_{2}}-\frac{\mathbb{I}_{n}}{\sqrt{n}}\right\|_{2}=\sqrt{2}$. This indicates that the $p$-th column of the output matrix $f(\mathbf{X}^{\star})$ is not parallel to $\mathbf{I}_{n}$ for any $p$. This further indicates that the output matrix does not have the identical vector in each row. ∎
@@ A.4 Technical Lemma
To demonstrate (23), we equivalently show
@@
Clearly, (27) holds true for any $x\in(0,1)$. ∎
@@ Appendix B Experiment Details
To more intuitively compare the security differences between the SOLID method and a fully-secured approach, we define $\Delta\textbf{ADR}(I)=\text{ADR}(I)-\text{ADR}([L])$ to assess the resilience of the secured set $I$ relative to the fully-secured approach. A smaller value of $\Delta\textbf{ADR}$ indicates resilience similar to that of the fully-secured model.
@@
The foundation models we use in our experiments are selected from open-source repositories, and Table 3 shows the basic information of the models and their sources. Specifically, we employ Llama2-70B-chat [^1] https://huggingface.co/meta-llama/Llama-2-70b-chat-hf, Llama2-7B-chat [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, and Mistral-7B-v0.1333https://huggingface.co/mistralai/Mistral-7B-v0.1. For smaller models, we select Phi-2444https://huggingface.co/microsoft/phi-2 and Phi-1.5555https://huggingface.co/microsoft/phi-1_5. We also consider OPT model [^6] https://huggingface.co/facebook/opt-350m, which has only 350 million parameters and 24 decoder layers.
@@
B.2 Distillation Attacks.
Attack implementation details. In performing FT-all and FT-secure model distillation attacks, we adhere to the training hyper-parameters outlined in the Llama2 report (Touvron et al., 2023), employing the AdamW optimizer with a cosine learning rate scheduler. The initial learning rate is set to $2\times 10^{-5}$, with a weight decay of 0.1, a batch size of 128, and bfloat16 precision for input sequences of 512 tokens. The LLaMA2-70B model is trained for 3 epochs with a random seed of 42, while other models are trained for 5 epochs across three seeds: 42, 1234, and 20. Despite limiting training to 3 epochs for the 70B model, the training loss stabilized effectively. Our implementation builds upon the llama-recipes repository provided by META.
For SEM attacks, distinct configurations were employed for SOLID and SAP-DP. In the case of SOLID, hidden representations from the secure-source components were collected and paired with the input data to train a substitute model. In contrast, for SAP-DP, representations from the sixth decoder layer and the model’s final logits were utilized to construct the training dataset. In accordance with (Tamber et al., 2024), we applied a learning rate of 1.5e-4, a weight decay of 0.01, and a linear learning rate scheduler with 500 warmup steps. Both training and validation batch sizes were set to 32, with MSE as the loss function. SOLID was trained for 30 epochs due to its smaller model size, whereas SAP-DP was trained for 5 epochs.
All distillation experiments were conducted on Nvidia 4090 24G, 6000 Ada 48G, and A100 80G GPUs, utilizing PyTorch 2.2.0 and CUDA 11.8 on Ubuntu 20.04.6 LTS.
Base 51k Distillation Dataset. We ensure dataset coverage and reliability by using a 1:1 ratio of the MMLU auxiliary training set  [^7] https://github.com/hendrycks/test and Alpaca dataset  [^8] https://github.com/tatsu-lab/stanford_alpaca/blob/main/alpaca_data.json, extracting 25.5k samples from each. From the MMLU auxiliary training data (Hendrycks et al., 2021), we sample 50%, and from Alpaca (Taori et al., 2023), we use a step size of 2 to enhance diversity. The datasets are then formatted for model training, applying Alpaca and MMLU prompts from Table 4.
@@
Validation Datasets. Table 5 outlines the composition of the validation dataset. We extracted 50% from each of the 57 MMLU validation sub-datasets, totaling 1.5K instances, paired with Alpaca data selected using a step size of 751.
@@
Total Length 1516 1516
@@
In this section, we provide further details on the baselines used in our comparisons: SAP-DP and fully-secured. These schemes represent different strategies, each with distinct trade-offs in terms of customizability and security against model distillation attacks.
SAP. The Split-and-Privatize (SAP) framework (Shen et al., 2023) offers an approach to balance between protecting model privacy and data privacy while maintaining competitive performance. Specifically, the SAP framework keeps the bottom six encoder layers open, allowing user access and fine-tuning while securing the deeper layers on the vendor.
@@
Fully-secured. Following (Eiras et al., 2024), we use the fully-secured approach as a baseline. This assumes the adversary has no access to internal model parameters, treating the model as a black-box, where only output data can be collected. We slightly broaden this setup by assuming the adversary knows the model’s architecture but no other details. Thus, distilling the fully-secured model involves using the collected data to retrain a model with the same architecture to restore its general functionality.
DarkneTZ. Based on the work of (Mo et al., 2020), we use DarkneTZ as a baseline to test whether protecting only the output layers is sufficient to defend against distillation attacks. In this setup, we assume the adversary has no access to the model parameters of the output layers, specifically the last decoder layer. Similar to the SAP framework, this approach allows the adversary to access and fine-tune all layers except the final decoder layer.
B.4 Implementation Details of SOLID.
Evaluation Datasets. We created a 1.5K Evaluation Set to assess model security under various secure-sourcing strategies. This set includes 50% of entries from each of the 57 MMLU validation sub-datasets (Hendrycks et al., 2021), distinct from Validation Set outlined in Table 5. Additionally, we selected an equal number of Alpaca dataset (Taori et al., 2023), using a step size of 751, ensuring no overlap with the Validation Set.
Hyper-parameter Sensitivity. As shown in Figure 6, we evaluate SOLID’s sensitivity to tolerance magnitude $\varepsilon$, adjusting it from 0.05 to 1 in 0.05 increments while calculating the $\Delta\text{ADR}$ for six distilled models. The results indicate that SOLID is minimally sensitive to changes in $\varepsilon$, with $\Delta\text{ADR}$ values stabilizing as $\varepsilon$ increases. This stability arises from the need for a smaller secured layer at higher $\varepsilon$, allowing the condition $R(I)\leq(1+\varepsilon)R([L])$ to be met with fewer layers. Additionally, the increase in $\Delta\text{ADR}$ is smaller for larger models, suggesting that privatizing more parameters beyond a certain point offers diminishing returns in security.